OpenStudy (anonymous):

If 18 is subtracted from twice the square of an integer, the result is equal to nine times the integer. Find the integer.

OpenStudy (anonymous):

Thank you but I would like to know whereyou get the -3/2 and 6 from @pooja195

OpenStudy (anonymous):

let x = the integer: 2x^2 - 18 = 9x 2x^2 - 9x - 18 = 0 Now either factor that or use the quadratic formula if you're lazy like me. x = -3/2 and 6, so x must be 6 because it has to be an integer. Done!